# Interpreting token compositionality in LLMs: A robustness analysis

### v1

@@ Abstract
Understanding the internal mechanisms of large language models (LLMs) is integral to enhancing their reliability, interpretability, and inference processes. We present Constituent-Aware Pooling (CAP), a methodology designed to analyse how LLMs process compositional linguistic structures. Grounded in principles of compositionality, mechanistic interpretability, and information gain theory, CAP systematically intervenes in model activations through constituent-based pooling at various model levels. Our experiments on inverse definition modelling, hypernym and synonym prediction reveal critical insights into transformers’ limitations in handling compositional abstractions. No specific layer integrates tokens into unified semantic representations based on their constituent parts. We observe fragmented information processing, which intensifies with model size, suggesting that larger models struggle more with these interventions and exhibit greater information dispersion. This fragmentation likely stems from transformers’ training objectives and architectural design, preventing systematic and cohesive representations. Our findings highlight fundamental limitations in current transformer architectures regarding compositional semantics processing and model interpretability, underscoring the critical need for novel approaches in LLM design to address these challenges.
@@ 1 Introduction
The number and diversity of applications for Transformer-based large language models (LLMs) are currently at an accentuated growth, considering their performance at major NLP tasks continues to exceed expectations. Yet, characteristics linked to LLMs critical deficiencies, such as hallucinations and lack of explainability, remain poorly elucidated despite intense research efforts. In particular, there is the matter of linguistic compositionality: how different units of text (morphemes, words, phrases) are combined into units of meaning, and how this can related to models of language interpretation.
Figure 1: Illustration of the CAP process. Constituent Segmentation creates segments based on token constituents, representing linguistic units like words or phrases. CAP is applied at layer $m$, pooling activations using aggregation protocols (e.g., max, mean, sum). CAP can be applied to any component, provided dimensionality is maintained. Activations at layer $m+1$ show the model’s new-reduced dimensionality. The results graph reflects how CAP is applied at different depths, with accuracy recorded for comparison..
Uncovering compositional representations in Transformer-based models is a key step in reducing the semantic gap between user intent and model interpretation, explaining model behaviour in different settings and mitigating hallucinations and inconsistent reasoning. To this effect, current work has predominantly focused on mapping input-outputs pairs Yin et al. (2024), input embeddings Haslett and Cai (2024) and layer-wise outputs Yu and Ettinger (2020); Modarressi et al. (2023) to expected latent representations by means of semantic similarity or correlation. Such focus reflects two intuitive expectations regarding LLM behaviours:
@@
Prior work has shown, through different perturbation experiments, that LLM representations and responses are fragile and unreliable across models Wang et al. (2023); Fodor et al. (2024); Hu et al. (2024). Furthermore, analysis of phrasal representations Yu and Ettinger (2020); Carvalho et al. (2022) revealed no significant correlation with phrasal compositional semantics. However, the mechanisms behind such obfuscation of the knowledge addressable by a Transformer model are still open research, which motivates this work.
Aiming to elucidate how token representations, in particular w.r.t. compositionality, affect LM interpretation, this work investigates the sensitivity/robustness of LLMs across local compositional perturbations at different stages of a Transformer model. In this context, we propose Constituent-Aware Pooling (CAP), a perturbation method for pooling LLM activations that correspond to individual tokens into activations representing more coherent linguistic units, using it to improve the understanding of compositionality mechanisms within transformer-based LMs (for an overview of compositionality and localisation, see Appendix A).
Through this intervention, contrary to expectations of incremental compositional/semantic buildup, we found that LLMs focus heavily on isolated token features rather than hierarchically integrating semantic information across layers, leading to significant performance drops when compositional activation perturbations (CAP) are applied. These fluctuations, especially pronounced in early and middle layers, indicate that aggregation of syntactic and semantic information is distributed across multiple layers rather than localised in any single layer. Surprisingly, larger models are more fragile to compositional perturbations than smaller ones, highlighting the significant difference in the way Transformer-based LMs build meaning representations.
@@
Based on the empirical results and theoretical model presented here, we postulate that traditional compositional semantic representations do not exist in any particular stage of a standard Transformer model. This is independent of model size, supervision type or inference task, but linked to the number of hidden layers. Our findings point toward the use of specialised architectures and/or training objectives in order to elicit such representations. The supporting datasets and software are available at anonymous url.
@@ 2 Tokenisation and compositionality in LLMs
The observed disconnection between LLM internal representations and linguistic knowledge regarding compositionality raises practical and theoretical questions towards the robustness of such models to perturbations strictly tied to compositional semantics (Appendix A). Such questions are specially relevant in solving semantic gaps between input prompts and expected responses, as well as localising linguistic knowledge and improving explainability. One way in which they can be addressed is by systematically assessing the impact of said perturbations on model inference performance, at each model layer. We elaborate the methodology to achieve this goal in the following section.
@@ 3 Assessing compositional aggregation robustness
In order to accurately measure the effects of compositional aggregation at different layers of abstraction within the model, the inference objective should be a task that is both: 1) strictly dependent on the input tokens and their composition, with few possible input variations; 2) contain as few tokens as possible in the output. For this reason, the following tasks were selected (Figure 1):
@@
The reader is referred to Appendix B.1 for a full formal definition of the tasks and their associated input formats.
Regarding a method to deliver compositional perturbations, we propose Constituent-Aware Pooling (CAP), an approach for pooling LLM activations that correspond to individual tokens into activations representing cohesive linguistic units. CAP operates at two levels: (i) word-level: Grouping tokens that form a single word, and (ii) phrase-level: Grouping tokens that form a single phrase.
For the word-level, CAP reverse-maps each model’s tokeniser to reconstruct complete words and determine their ranges. At the phrase-level, it employs a syntactic parser (e.g., Benepar parser Kitaev et al. (2019); Kitaev and Klein (2018)) to match tokens to corresponding phrasal constituents and establish their ranges. Aggregations are done progressively from the first to the last layers, using three different pooling modes: Max aggregation, which takes the maximum activation within the segment; Mean aggregation, which averages the activations across the segment; and Sum aggregation, which sums the activations across the segment, capturing cumulative information loss. These aggregation types (max, mean, and sum) offer different perspectives on how information is processed and retained, with Max aggregation highlighting critical features, while Mean and Sum provide broader views of overall performance sensitivity and cumulative loss across tokens.
@@
Formally, we focus on a mathematical conceptualisation of transformers [^1] In this paper, the term ’transformer models’ specifically refers to autoregressive transformers. introduced by Elhage et al. (2021). In Transformer architectures, computation is organised into sequential layers or residual blocks that interact via a shared data pathway (residual stream). Each layer reads input from the residual stream, processes it through its components, and writes the output back to the residual stream. The primary components within these layers are attention heads and feed-forward (FF) neural networks. Attention heads are responsible for transferring information between tokens through the self-attention mechanism, allowing each token to attend to others in the sequence. Feed-forward layers apply non-linear transformations independently to each token representation, enhancing the model’s expressive capacity. The residual stream serves to store and propagate information across layers, enabling the integration of new outputs with existing representations while preserving original input information through residual connections. The structure of transformers provides a basis for processing activations within these model components, enabling the systematic evaluation of compositionality at various layers.
Given a model with $L$ layers and an input sequence of length $K$, let ${(s_{i},e_{i})}$ represent the set of multiple syntactic units ranges, where each $(s_{i},e_{i})$ corresponds to a word, a phrase or other linguistic constituents. We apply CAP over these unit ranges to key model components, including attention heads, residual streams, and FF layers, using a pooling function $\alpha$. CAP results in reducing the dimension corresponding to input tokens from $K$ to $K-\Sigma_{i}(e_{i}-s_{i})$. Below, we formalise how CAP is applied for different components. For simplicity, we demonstrate the operation over one unit, but this can be extended to any number of units. It can also be applied to any component, provided that the dimensional requirements for information flow, as described in Elhage et al. (2021), are respected.
Attention Heads. Let $A\in\mathbb{R}^{B\times H\times K\times K}$ be the attention activation, where $B$ is the batch size, $H$ is the number of attention heads, and $K$ is the sequence length, we apply CAP over the token range $(s_{i},e_{i})$ for a syntactic constituent as follows:
$\begin{split}A_{\text{grouped}}[b,h,i,j]&=\alpha_{t=s_{i}}^{e_{i}}\alpha_{t^{\prime}=s_{j}}^{e_{j}}A[b,h,t,t^{\prime}]\\ A&=A_{\text{grouped}}\end{split}$ (1)
where $\alpha$ represents the pooling function applied over the specified token ranges $(s_{i},e_{i})$ and $(s_{j},e_{j})$, smoothing attention scores across tokens within the same syntactic constituent.
Feed-Forward (FF) Layers. Let $P\in\mathbb{R}^{B\times K\times H}$ denote the FF activations, where $B$ is the batch size, $K$ is the sequence length, and $H$ is the FF dimension. CAP is applied the over the constituent range $(s_{i},e_{i})$ as follows:
$\begin{split}P_{\text{grouped}}[b,i]&=\alpha_{t=s_{i}}^{e_{i}}P[b,t]\\ P&=P_{\text{grouped}}[b,i]\end{split}$ (2)
Residual Stream. Let $R\in\mathbb{R}^{B\times K\times H}$ represent the residual stream activations, where $B$ is the batch size, $K$ is the sequence length, and $H$ is the hidden dimension, we apply CAP over the range $(s_{i},e_{i})$ as follows:
$\begin{split}R_{\text{grouped}}[b,i]&=\alpha_{t=s_{i}}^{e_{i}}R[b,t]\\ R&=R_{\text{grouped}}\end{split}$ (3)
Here, $\alpha$ serves as the aggregation function, pooling the activations within the token range $(s_{i},e_{i})$ into a single representation.
These direct interventions within the model are then evaluated with regard to a downstream accuracy effect by measuring an accuracy difference, $\Delta A=A_{o}-A_{c}$, where $A_{c}$ is post-CAP accuracy. As interventions can be localised within the $A,P,R$ components, layer-wise, CAP allows for a granular characterisation of the compositional behaviour across the model structure. In broader terms, it is assumed that a more localised compositional behaviour across layers would imply in a lower $\Delta$.
@@
Figure 2: Average grouped accuracy of CAP across different aggregation functions for normalised layer positions (0%-100%) is shown for word-level CAP (TW) and phrasal-level CAP (TP). The shaded areas represent the variance in model performance, with larger shaded regions indicating greater variability across layers. Sub-figures (a)-(c) illustrate the CAP effect on the original (Org) models, while sub-figures (d)-(f) show its impact on the fine-tuned (FT) models. Fine-tuning consistently improves performance, particularly in the middle to late layers (25%-100%), while early layers (0%-25%) show more variability and lower accuracy across models.
@@ 4.1 Experimental setup & datasets
Metrics.
To evaluate model performance, we report original accuracy ($A_{o}$) and accuracy drop after grouping ($\Delta A=A_{o}-A_{c}$, where $A_{c}$ is post-CAP accuracy). These metrics effectively capture the models’ performance and the impact of CAP. $\Delta A$ directly quantifies the grouping effect, enabling clear cross-model and cross-task comparisons. An expanded reporting on the results is in Appendix C.2.
@@
Datasets
Experiments were conducted on three WordNet-derived datasets: definitions, hypernyms, and synonyms Fellbaum (1998). These correspond to the three tasks: IDM, HP, and SP. Initially, models were tested on these constructed datasets using the correctly predicted samples as the baseline for further analysis (see Appendix B.2 for datasets details).
@@
The methodology was tested across various decoder-only transformer models Vaswani (2017). Our main focus was on GPT-2 (small: 124M, medium: 355M, large: 774M parameters) Radford et al. (2019), Gemma (2B parameters) Team et al. (2024), and Llama-3 (8B parameters) Dubey et al. (2024). The models differ in tokenisation methods: GPT-2 models use byte-level BPE tokenisation, Llama-3 employs byte-level BPE but with a large 128,000-token vocabulary, while Gemma-2B utilises a SentencePiece-based Kudo and Richardson (2018) tokeniser. All models were tested with and without fine-tuning on downstream tasks. Fine-tuning was conducted for three epochs with a learning rate of 5e-5. The models were chosen to represent a range of sizes, architectures, and tokenisation strategies (see Appendix B.3 for further details on the models and fine-tuning parameters).
@@ 4.2 Results and discussion
Compositional inference in LLMs is not a purely incremental process. Contrary to expectations of a smooth and steady layer-wise performance improvement, we observe fluctuations when CAP is applied across layers (see Figure 2). Under perturbations, a more significant drop in performance occurs in the early and middle layers, followed by a sharp improvement (Figure 2 (a), (c), (e), and (f)), suggesting that these layers struggle to process the CAPed activations, particularly the linguistic features captured in earlier layers. Instead of progressively and hierarchically building up semantic information from individual tokens to more complex phrases and words, the models appear to focus heavily on isolated token features. The results indicate that attention is distributed over the input tokens and model layers in a non-systematic and decentralised manner that is highly context-dependent and shows minimal reliance on the sequential or positional relationships of constituents. This phenomenon is particularly evident in the sharp decline in performance in the SP and HP tasks, where contextual information is limited and while applying CAP on a phrasal level, effectively abstracting the contextual information. We argue that this behaviour results from the model’s training objective, which seeks to maximise information gain from each layer towards the predicted tokens, at the cost of reducing mutual information between tokens in a single layer. This means that aggregation, including syntactic, is performed across multiple layers and thus is not localisable from any single given layer. An information theoretical analysis elaborates this reasoning in Section 5. Our findings highlight the sensitivity of token representations to compositional structures, suggesting that the performance fluctuations can be attributed to information loss incurred as a function of token mutual information across layers.
Larger models are more fragile to compositional perturbations. Larger models are more sensitive to early-layer interventions than smaller models. For instance, while applying word-level CAP, GPT2-large maintains an accuracy close to 90% in both the original and fine-tuned cases for the IDM task, while Llama3-8B shows an early-layer accuracy of around 75% in both cases (Figure 2 and Table 1). Gemma-2B experiences greater variability, with a sharp accuracy drop of 85%-90% in the early layers across most tasks and settings, highlighting its sensitivity to perturbations. In contrast, smaller models like GPT2-small and GPT2-medium perform more steadily, with early layers’ accuracy dropping by around 10-20%. While larger models typically outperform smaller ones in a range of tasks (see Appendix C.1 for original models’ performance without intervention), they exhibit a greater reliance on the identification of intrinsic features in the early layers. We also observe that the performance of larger models, Gemma-2B and Llama3-8B, is completely broken when applying phrasal-level CAP. This finding suggests a stronger dependency on layer-wise information gain in larger models, where they attempt to separate features in an uncorrelated and highly distinct manner within their parameters. While this ability gives them an advantage in identifying complex and unique feature patterns, it also renders them more vulnerable to changes or noise in the context—a weakness that threatens the models’ robustness and integrity.
@@
Activation abstraction vs the information loss. For the word-level CAP, Table 1 shows the accuracy drop over a sample of models for the IDM task and the different aggregation functions, presented in a relative manner as to layers positions (see Appendix C.2 for the rest of models and tasks results). The results indicate significant disparities among the aggregation functions’ performance, most notably with the Max aggregation. Larger models, such as Gemma-2B and Llama3-8B, were the most affected by this aggregation type. This finding supports our argument that these models tend to distribute information in a fragmented manner, lacking the integration of compositional (lexical and semantic) information across tokens and contiguous layers. Thus, they tend to exhibit high sensitivity to any perturbation in their activations. The Mean aggregation provides a more balanced method, consistently resulting in more moderate accuracy drops. However, even with this balanced aggregation, the drop in performance suggests that there is no identifiable consistent and robust mechanism for composing token meanings. This issue is more prominent in the token-phrases experiments, as shown in Figure 2. Interestingly, the Sum aggregation performed better in fine-tuned models, while the Mean aggregation was slightly more effective in the original models. Nevertheless, the differences between them were relatively small. The Sum aggregation reflects the cumulative effect of aggregating tokens into larger segments, reinforcing our earlier conclusion. Instead of progressively building semantic information across layers, the models exhibit a cumulative loss of information, particularly when interventions occur in the early layers.
Fine-tuning expedites the recovery process. As seen in Figure 2 (d-f) and Table 1, there is a positive correlation between the models’ ability to maintain performance and fine-tuning (FT). The biggest improvements are observed in the 75%-100% layer positions. SP demonstrated the best gain from FT, which we attribute to the task’s high specificity and the minimal activation reduction under CAP. Max aggregation shows the greatest improvement with FT, likely due to fine-tuning amplifying dominant activations and helping retain key information. For example, in Gemma-2B, the drop decreased from 97.91% to 57.65% in the 1% layer, and in Llama3-8B, from 14.98% to around 7% in the 75% layer. Mean aggregation also benefited from FT, with the accuracy drop in Gemma-2 B’s 75% layer reduced from 31.03% to 15.00%. However, except for Gemma-2B, Sum aggregation saw only slight improvements. For smaller models like GPT2-large, results showed either the opposite behaviour or minimal improvement, which we believe is due to overfitting. While fine-tuning speeds up recovery, it does not fully mitigate performance reduction, indicating that current Transformer architectures still struggle to form compositional semantic representations from individual tokens.

### v2

@@ Abstract
Understanding the internal mechanisms of large language models (LLMs) is integral to enhancing their reliability, interpretability, and inference processes. We present Constituent-Aware Pooling (CAP), a methodology designed to analyse how LLMs process compositional linguistic structures. Grounded in principles of compositionality, mechanistic interpretability, and information theory, CAP systematically intervenes in model activations through constituent-based pooling at various model levels. Our experiments on inverse definition modelling, hypernym and synonym prediction reveal critical insights into transformers’ limitations in handling compositional abstractions. No specific layer integrates tokens into unified semantic representations based on their constituent parts. We observe fragmented information processing, which intensifies with model size, suggesting that larger models struggle more with these interventions and exhibit greater information dispersion. This fragmentation likely stems from transformers’ training objectives and architectural design, preventing systematic and cohesive representations. Our findings highlight fundamental limitations in current transformer architectures regarding compositional semantics processing and model interpretability, underscoring the critical need for novel approaches in LLM design to address these challenges.
@@ 1 Introduction
The number and diversity of applications for Transformer-based large language models (LLMs) are currently at an accentuated growth, considering their performance at major NLP tasks continues to increase. Yet, the understanding of characteristics linked to LLMs critical deficiencies, such as hallucinations and lack of interpretability, remain limited. In particular, there is the matter of linguistic compositionality: how different units of text (morphemes, words, phrases) are combined into units of meaning, and how this relates to models of language interpretation.
Figure 1: Illustration of the CAP process. Constituent Segmentation creates segments based on token constituents, representing linguistic units like words or phrases. CAP is applied at layer $m$, pooling activations using aggregation protocols (e.g., max, mean, sum). CAP can be applied to any component, provided dimensionality is maintained. Activations at layer $m+1$ show the model’s new-reduced dimensionality. The results graph reflects how CAP is applied at different depths, with accuracy recorded for comparison.
Uncovering compositional representations in Transformer-based models is a key step in reducing the semantic gap between user intent and model interpretation, explaining model behaviour in different settings and mitigating hallucinations and inconsistent reasoning. To this effect, current work has predominantly focused on mapping input-output pairs Yin et al. (2024), input embeddings Haslett (2024) and layer-wise outputs Yu and Ettinger (2020); Modarressi et al. (2023) to expected latent representations by means of semantic similarity or correlation. Such focus reflects two intuitive expectations regarding LLM behaviour:
@@
Prior work has shown, through different perturbation experiments, that LLM representations and responses are fragile and unreliable across models Wang et al. (2023); Fodor et al. (2024); Hu et al. (2024). Furthermore, analysis of phrasal representations Yu and Ettinger (2020); Carvalho et al. (2025) revealed no significant correlation with phrasal compositional semantics. However, the mechanisms behind such obfuscation of the knowledge addressable by a Transformer model are still open research questions, which motivates this work.
Aiming to elucidate how token representations, particularly with respect to compositionality, affect LLM interpretation, this work investigates the sensitivity/robustness of LLMs across local compositional perturbations at different stages of a Transformer model. In this study, we specifically focus on autoregressive decoder Transformers, investigating their compositional properties. We propose Constituent-Aware Pooling (CAP), a perturbation method for pooling LLM activations that correspond to individual tokens into activations representing more coherent linguistic units, using it to improve the understanding of compositionality mechanisms within Transformer-based LMs (for an overview of compositionality and localisation, see Appendix A).
Through this intervention, contrary to expectations of incremental compositional/semantic build-up, we found that LLMs focus heavily on isolated token features rather than hierarchically integrating semantic information across layers, leading to significant performance drops when CAP is applied. These fluctuations, especially pronounced in early and middle layers, indicate that aggregation of syntactic and semantic information is distributed across multiple layers rather than localised in any single layer. Surprisingly, larger models are more fragile to compositional perturbations than smaller ones, highlighting the significant difference in the way Transformer-based LMs build meaning representations.
@@
Based on the empirical results and theoretical model presented here, we postulate that traditional compositional semantic representations cannot be isolated to any particular (intermediate) stage of a standard Transformer model. This is independent of model size, supervision type or inference task, but linked to the number of hidden layers. Our findings point toward the use of specialised architectures and/or training objectives in order to elicit such representations. The supporting datasets and software are available at a public repository111¡anonymised url¿.
@@ 2 Tokenisation and compositionality in LLMs
The observed disconnection between LLM internal representations and linguistic knowledge regarding compositionality raises practical and theoretical questions towards the robustness of such models to perturbations strictly tied to compositional semantics (Appendix A). Such questions are especially relevant in solving semantic gaps between input prompts and expected responses, as well as localising linguistic knowledge and improving interpretability. One way in which they can be addressed is by systematically assessing the impact of said perturbations on model inference performance, at each model layer. We elaborate on the methodology to achieve this goal in the following section.
@@ 3 Assessing compositional aggregation robustness
To accurately assess the effects of compositional grouping at different layers of abstraction within transformer models, the inference objective should be a task that is both: 1) strictly dependent on the input tokens and their composition, with few possible input variations; 2) contain as few tokens as possible in the output. For this reason, the following tasks were selected (Figure 1):
@@
Formal task definitions and input formats are detailed in Appendix B.1.
Constituent-Aware Pooling (CAP). To introduce compositional perturbations, we propose CAP, a method for pooling (i.e., grouping) LLM activations corresponding to individual tokens into cohesive linguistic units. CAP operates at two levels: (i) word-level: grouping tokens that form a single word, and (ii) phrase-level: grouping tokens that form a single phrase. At the word-level, CAP reverse-maps each model’s tokeniser to reconstruct complete words and identify their activation ranges. At the phrase-level, CAP uses a syntactic parser, such as Benepar Kitaev et al. (2019); Kitaev and Klein (2018), to align tokens with their corresponding phrasal constituents and define their activation ranges. Further details on the parser evaluation methodology are provided in Appendix D.
CAP Pooling Protocols. CAP is applied progressively across layers using three protocols $\alpha$: Max: selects the maximum activation within a segment, identifying dominant features and their propagation through layers; Mean: computes the average activation, providing a balanced representation of all token contributions and their collective impact on model decisions; and Sum: sums the activations, capturing cumulative information flow and aggregates effects of token interactions. These protocols offer complementary insights into how models process and integrate information: Max reveals feature prominence patterns, Mean shows distributed representation effects, and Sum reflects accumulated semantic content across segments.
@@
Transformer conceptualisation and the formalisation of CAP. This work builds on the mathematical framework of transformers introduced by Elhage et al. (2021), where computation is formalised into sequential residual blocks. Each layer reads inputs from the residual stream, processes them through its components (attention heads and feed-forward neural networks (FF)), and writes the outputs back into the residual stream. Attention heads are responsible for transferring information between tokens through the self-attention mechanism, allowing each token to attend to others in the sequence. FF apply non-linear transformations independently to each token representation, enhancing the model’s expressive capacity. The residual stream stores and propagates information across layers, enabling the integration of new outputs with existing representations while preserving original input information through residual connections. Let the transformer model have $L$ layers, input sequence of length $K$, batch size $B$, and inner activations $X$, with with tensor shapes varying by model component as follows:
  • Attention layers output: $X\in R^{B\times K\times H_{m}}$, where $H_{m}$ is the hidden dimension after projection.
  • FF: $X\in R^{B\times K\times H_{f}}$, where $H_{f}$ is the feed-forward dimension.
  • Residual stream: $X\in R^{B\times K\times H_{h}}$, where $H_{h}$ is the hidden dimension.
Let $\mathcal{S}=\{(s_{1},e_{1}),\dots(s_{n},e_{n})\}$ be the set of syntactic unit ranges (e.g., tokens, words or phrases), where $s_{i}$ and $e_{i}$ denote the start and end indices of the $i$-th range. CAP pools/groups activations within these ranges, reducing the sequence dimension $K$ to a grouped dimension $G$, where
$G=K-\Sigma_{i=1}^{n}(e_{i}-s_{i})$ (1)
For each syntactic unit, CAP applies the grouping function $\alpha$ over the range $[s_{i},e_{i}]$ in one of three ways, formalised as follows:
[…]
The grouped activations transform as follows:
  • For attention layers output, $X\in R^{B\times K\times H_{m}}$ becomes $X\in R^{B\times G\times H_{m}}$.
  • For FF, $X\in R^{B\times K\times H_{f}}$ becomes $X\in R^{B\times G\times H_{f}}$.
  • For residual stream:, $X\in R^{B\times K\times H_{h}}$ becomes $X\in R^{B\times G\times H_{h}}$.
This process consolidates activations for each syntactic unit, enabling systematic evaluation of compositional robustness across layers. For simplicity, we demonstrate the operation over these components, but this approach can be extended to any transformer’s components, provided that the dimensional requirements for information flow, as described in Elhage et al. (2021), are respected. For example, consider attention layer internal activations of shape $X\in\mathbb{R}^{B\times H_{a}\times K\times K}$, where $H_{a}$ is the number of attention heads, and $K$ represents the query and key token dimensions. Applying CAP with the Sum protocol involves aggregating activations over the query range $[s_{i},e_{i}]$ and the key range $[s_{j},e_{j}]$. The grouped activations are computed as: $\alpha([s_{i},e_{i}],[s_{j},e_{j}])=\sum_{t=s_{i}}^{e_{i}}\sum_{t^{\prime}=s_{j}}^{e_{j}}X[b,h,t,t^{\prime}]$. After applying CAP, the grouped activations have the shape $X\in\mathbb{R}^{B\times H_{a}\times G\times G}$, where $G$ is the number of grouped syntactic units. This ensures that query-key interactions are consolidated into cohesive syntactic units, aligning activations with higher-level linguistic structures. We examine CAP’s reduction ratio ($K\to G$) at the word-level and its effects across models, with detailed analysis in Appendix C.
The CAP effect on models is evaluated by measuring their accuracy post-CAP on a baseline test consisting of examples correctly predicted by the original models. This ensures that the evaluation focuses on instances where CAP directly tests compositional robustness. Specifically, we report three key metrics: the original accuracy ($A_{o}$), which represents the model’s accuracy on the baseline test before applying CAP and establishes a reference for evaluating the grouping effect; the grouped accuracy ($A_{c}$), which measures the model’s accuracy post-CAP, averaged across all CAP protocols (sum, mean, max) and reflects how well the model retains its predictions after compositional grouping; and the accuracy drop ($\Delta A$), defined as $\Delta A=A_{o}-A_{c}$, which quantifies the performance loss due to CAP, where lower $\Delta A$ values indicate more robust compositional behaviour and better preservation of semantic information across layers. These metrics offer a framework for comparing tasks and models, allowing a granular assessment of compositional representations.
@@
Figure 2: Average grouped accuracy of CAP across different aggregation functions for normalised layer positions (0%-100%) is shown for word-level CAP (TW) and phrasal-level CAP (TP). Sub-figures (a)-(c) illustrate the CAP effect on the original (Org) models, while sub-figures (d)-(f) show its impact on the fine-tuned (FT) models. Fine-tuning consistently improves performance, particularly in the middle to late layers (25%-100%), while early layers (0%-25%) show more variability and lower accuracy across models.
@@ 4.1 Experimental setup & datasets
Datasets and metrics.
The CAP effect is evaluated using three WordNet-derived datasets—definitions, hypernyms, and synonyms—corresponding to the IDM, HP, and SP tasks Fellbaum (1998). Test examples correctly predicted by the original models ($A_{o}$) form the baseline for subsequent CAP evaluation. Grouped accuracy ($A_{c}$) is calculated post-CAP for this subset, ensuring that CAP’s effect is isolated to examples where the original models performed correctly. The drop in accuracy ($\Delta A$) is reported per protocol (sum, mean, max) to assess the impact of different aggregation methods on model performance. See Appendix B.2 for dataset details and Appendix E.2 for comprehensive results.
@@
The methodology was tested across various decoder-only transformer models Vaswani (2017). Our main focus was on GPT-2 (small: 124M, medium: 355M, large: 774M parameters) Radford et al. (2019), Gemma (2B parameters) Team et al. (2024), Llama (3B, 8B, and 8B-instruct parameters) Dubey et al. (2024), and Qwen (0.5B, 1.5B, and 3B parameters) Yang et al. (2024). These models use different tokenisation approaches: byte-level BPE (GPT-2, Qwen), expanded BPE with 128K vocabulary (Llama3), and SentencePiece (Gemma). Models were tested before and after task-specific fine-tuning (3 epochs, learning rate 5e-5). This selection spans diverse architectures, sizes, and tokenisation strategies (see Appendix B.3 for further details on the models and fine-tuning parameters).
@@ 4.2 Results and discussion
Compositional inference in LLMs is not a purely incremental process. Contrary to expectations of a smooth and steady layer-wise performance improvement, we observe significant fluctuations when CAP is applied across layers (Figure 2). Performance drops notably in early and middle layers, followed by sharp improvements (Figure 2 (a)-(c), (e), and (f)), suggesting these layers struggle to process CAPed activations, particularly the pooled linguistic features captured in earlier layers. Rather than progressively building semantic information from individual tokens to complex phrases, the models appear to focus heavily on isolated token features. The results indicate that attention is distributed over input tokens and model layers in a non-systematic and decentralised manner that is highly context-dependent, showing minimal reliance on sequential or positional relationships of constituents. This phenomenon is particularly evident in the sharp decline in SP and HP tasks, where contextual information is limited during phrase-level CAP application. We argue that this behaviour stems from the model’s training objective, which maximises information gain in each layer towards predicted tokens at the cost of reducing mutual information between tokens in a single layer. This behaviour means that aggregation, including syntactic, is performed across multiple layers and thus is not localisable from any single given layer. An information theoretical analysis elaborates this reasoning in Section 5. Our findings highlight how compositional structures are highly sensitive to token representation dynamics across layers, suggesting that performance fluctuations can be attributed to information loss incurred as a function of token mutual information across layers.
Larger models are more fragile to compositional perturbations. The IDM task highlights this fragility in larger models, as larger models rely on finer feature extraction. Within families, distinct patterns emerge: original Qwen’s smaller variants show better IDM robustness (e.g., at position 25% there was a 7.69% drop on Qwen-1.5B vs 12.11% on Qwen-3B), while Llama3 exhibits capacity-dependent behaviour with the 3B variant being more vulnerable than 8B. Notably, Llama3-8B-Instruct, despite being fine-tuned, performed worse in IDM compared to Llama 8B and 3B, but excels in structured tasks (SP, HP), underscoring the role of architecture and training in compositional robustness. Despite having similar reduction ratios to Llama models (see Appendix C), Gemma-2B shows greater sensitivity to perturbations (e.g., at position 1% Max: Gemma-2B drops 97.91% vs. Llama3-8B’s 25.49%), likely due to its larger vocabulary enabling finer-grained tokenisation. While fine-grained token knowledge benefits standard tasks, it appears to increase susceptibility to compositional perturbations. The superior performance of Llama3-8B over its 3B variant can be attributed to its enhanced capacity for maintaining feature relationships across layers while preserving key compositional information. While larger models excel in standard tasks (see Appendix E.1), they exhibit a greater reliance on the identification of intrinsic features in the early layers. We find that phrasal-level CAP substantially impacts Gemma-2B and Llama models, suggesting a heavy dependence on layer-wise information gain, where they separate features in an uncorrelated and highly distinct manner. While this aids in identifying complex feature patterns, it also makes them more vulnerable to contextual noise—a weakness that threatens their robustness and integrity. Notably, Qwen models outperform Llama and Gemma despite similar parameter counts, likely due to byte-level BPE tokenisation and multilingual training, which enhance compositional stability, whereas Llama’s expanded BPE and Gemma’s SentencePiece prioritise efficiency over phrase retention, increasing vulnerability to CAP interventions.
@@
Activation abstraction vs the information loss. Table 1 reveals significant variations in aggregation function performance across sample models for the IDM task (see Appendix E.2 for the rest of the models and tasks results). The Max aggregation shows the most dramatic impact. This finding supports our argument that these models tend to distribute information in a fragmented manner, lacking the integration of compositional (lexical and semantic) information across tokens and contiguous layers. The Mean aggregation provides more balanced results, though performance drops still indicate absence of consistent compositional mechanisms. This issue becomes more pronounced in token-phrases experiments (Figure 2). The Sum aggregation consistently outperformed other methods, with Mean aggregation following closely behind, particularly in original models. The Sum aggregation reflects the cumulative effect of aggregating tokens into larger segments, reinforcing our earlier conclusion. Instead of progressively building semantic information across layers, the models exhibit cumulative information loss, particularly when interventions occur in early layers.
Fine-tuning enhances recovery capabilities across models. Figure 2 (d-f) demonstrates improved performance maintenance post-fine-tuning across all model families, with strongest gains in 75%-100% layer positions. SP tasks showed maximum benefit, attributed to high task specificity and minimal activation reduction under CAP. Max aggregation displayed the greatest improvement post-fine-tuning, likely due to enhanced retention of key information. For instance, Gemma-2B’s accuracy drop decreased from 97.91% to 57.65% in the 1% layer, while Qwen-3B improved from 7.23% to 3.65% in the 75% layer. Mean aggregation benefits were also substantial in smaller models, with Gemma-2B’s 75% layer drop reducing from 31.03% to 15.00%. The Qwen family showed consistent improvements across all aggregation types, though smaller models like GPT2-large demonstrated minimal gains, suggesting potential overfitting. Notably, larger models like Llama3-8B showed minimal gains from fine-tuning in IDM tasks, indicating that standard fine-tuning objectives may not directly enhance compositional robustness. Although fine-tuning strengthens models’ resilience under CAP, it does not fully resolve the challenge of forming stable compositional semantic representations, highlighting an architectural limitation in current transformer models.